Express the polynomial 1+3x+4x2 as a linear combination of the polynomials

1+2x+3x2, -1+x+x2 and 2+x+x2

1
Expert's answer
2021-08-12T18:13:55-0400

1+3x+4x2= a(1+2x+3x2) + b(-1+x+x2) + c(2+x+x2) = (a-b+2c) + (2a+b+c)x + (3a+b+c)x2

"\begin{cases}
 a-b+2c = 1 \\
 2a+b+c = 3 \\
 3a+b + c = 4
\end{cases}"

subtraction the second equation from the third gives: a =1

addition the first equation and the secong gives 3a + 3c = 4, and from this c = 1/3

from the fist equation: b = a +2c - 1= 2/3

so 1+3x+4x2= (1+2x+3x2) + 2/3 (-1+x+x2) + 1/3 (2+x+x2)